# LLaVE: Large Language and Vision Embedding Models with Hardness-Weighted Contrastive Learning

### v1

@@ 1 Introduction
Multimodal embedding models aim to encode inputs from any modality into vector representations, which then facilitate various multimodal tasks, such as image-text retrieval Wu et al. (2021); Zhang et al. (2024a), automatic evaluation Hessel et al. (2021), and retrieval-augmented generation (RAG) Zhao et al. (2023). Although advanced pretrained vision-language models such as CLIP Radford et al. (2021), ALIGN Jia et al. (2021), and SigLIP Zhai et al. (2023) can provide unified representations for text and images, they face difficulties when dealing with more complex tasks. Particularly, they adopt a dual-encoder architecture that encodes images and text separately, leading to poor performance in tasks such as interleaved image-text retrieval Wu et al. (2021).
In recent years, the rapid development and exceptional performance of large multimodal models (LMMs) have prompted researchers to focus increasingly on LMM-based multimodal embedding models. Compared to traditional pretrained vision-language models, LMMs not only demonstrate superior multimodal semantic understanding capabilities but also naturally support interleaved text and image inputs. This advantage makes them more flexible and efficient in handling multimodal embedding tasks. As a representative work, Jiang et al. (2024b) construct the Massive Multimodal Embedding Benchmark (MMEB), which encompasses 4 meta-tasks and 36 datasets, and train multimodal embedding models based on LMMs. Experiment results demonstrate that by providing suitable task instructions and employing contrastive learning for training, LMM can significantly outperform existing multimodal embedding models and generalize effectively across diverse tasks.
@@ 2 Preliminary Study
SUN397 ImageNet-R DocVQA TextVQA CIRR Wiki-SS-NQ MSCOCO RefCOCO
@@ 3 Our Framework
In this section, we first illustrate the inherent consistency between preference learning and contrastive learning. We then propose a simple yet effective framework, which mainly involves hardness-weighted contrastive learning (See Section 2) and cross-device negative sample gathering (See Section 3).
@@ 3.1 Hardness-Weighted Contrastive Learning
$\mathcal{L}_{i}=-log\frac{e^{{r_{\pi}(q_{i},t_{i})}}}{e^{{r_{\pi}(q_{i},t_{i})}}+\sum_{j\neq i}^{N}w_{ij}\cdot e^{{{r_{\pi}(q_{i},t_{j})}}}},$ (4)
where $w_{ij}$ represents the weight of learning difficulty. As shown in Figure 2, to estimate the learning difficulty of pairs, we introduce a reward model $r_{\theta}$ and set $w_{ij}=e^{r_{\theta}(q_{i},t_{j})}$. Accordingly, the policy model adjusts its learning of different negative pairs based on the feedback from the reward model. In this work, to achieve higher training efficiency and simpler implementation, we directly set the reward model $r_{\theta}$ to be the same as the policy model $r_{\pi}$, but without backpropagating gradients, i.e. $r_{\theta}(q_{i},t_{j})$ = $\alpha\cdot\mathrm{sg}(s_{ij})$, where $\alpha$ is a hyperparameter and $\mathrm{sg}(\cdot)$ denotes the stop-gradient operation. Note that $r_{\theta}$ can also adopt model structures other than the policy model. Finally, $\mathcal{L}_{i}$ is defined as
@@
From Equation 6, we can observe that the gradients of the negative pairs are proportional to the product $r_{\theta}(q_{i},t_{j})$, which implies that the greater the learning difficulty of a negative pair, the more significant its role in the gradient update.
@@ 3.2 Cross-Device Negative Sample Gathering
Figure 3: An example of cross-device negative sample gathering ($N$=$4$ and $K$=$3$). The plus signs represent positive pairs, and the minus signs represent negative pairs. Each device calculates the similarity between its own queries and the targets on all other devices, which is then used for loss computation.
@@ 4.1 Setup
Table 2: Results on the MMEB benchmark. IND represents the in-distribution dataset, and OOD represents the out-of-distribution dataset. In UniIR, the FF and SF subscripts under CLIP or BLIP represent feature-level fusion and score-level fusion, respectively. LLaVA-NeXT-7B-LR indicates the use of low-resolution (336×336) image inputs, while LLaVA-NeXT-7B-HR refers to the use of high-resolution (1344×1344) image inputs. The reported scores are the average Precision@1 over the corresponding datasets. The best results are marked in bold, and the second-best results are underlined. Part of the baseline results are sourced from Jiang et al. (2024b) and Zhou et al. (2024). Detailed results and qualitative evaluations can be found in Appendix A.3 and Appendix A.4.
@@
Our trained model, LLaVE, includes three scales: 0.5B, 2B, and 7B, based on LLaVA-OV-0.5B Li et al. (2024), Aquila-VL-2B Gu et al. (2024), and LLaVA-OV-7B Li et al. (2024), respectively. To facilitate community use, the training code for LLaVE is built on the widely-used Transformers Wolf et al. (2020) and DeepSpeed Rasley et al. (2020) packages. We use a batch size of 256, set the weighting hyperparameter $\alpha$ to 9  [^1] We empirically analyze the impact of $\alpha$ on model performance in Appendix A.1., and impose a total length limit of 4096. Furthermore, we employ the Higher Anyres technique Li et al. (2024) to support high-resolution images, setting the maximum image resolution to 672 $\times$ 672. The learning rate is set to 1e-5 for LLaVE-0.5B and LLaVE-2B, and 5e-6 for LLaVE-7B. For efficient training, we freeze the vision encoder and train the model for one epoch using the DeepSpeed ZeRO-3 strategy. Regarding training costs, LLaVE-0.5B and LLaVE-2B are trained on 8 NVIDIA A100 GPUs (40GB) for 12 and 17 hours, respectively, while LLaVE-7B is trained on 16 Ascend 910B GPUs (64GB) for 33 hours. More details can be found in Appendix A.2.
@@ 4.2 Main Results
Although previous models have achieved strong performance, our LLaVE series demonstrates consistent improvements over the best baseline across all metrics. LLaVA-NeXT-7B achieves a state-of-the-art overall score of 70.3, outperforming the previous SOTA model, MMRet, by 6.2 points and surpassing VLM2Vec (LLaVA-OV-7B) by 4.5 points. In grounding, LLaVE-7B attains an impressive score of 91.9, a +4.6 point improvement over VLM2Vec. LLaVE-7B also leads in VQA (65.4) and classification (65.7), with improvements of +4.3 and +2.2 points, respectively. In addition, we observe that the performance of LLaVE models scales consistently with model size, indicating that our framework has excellent scalability. It is worth mentioning that the performance of LLaVE-0.5B is already comparable to VLM2Vec (Phi-3.5-V-4B), while LLaVE-2B achieves an overall score of 65.2, surpassing the previously pretrained MMRet-7B that utilizes an additional 27M image-text pairs.
@@ 4.3 Ablation Study
Although the model (ID 6) already achieves near-SOTA performance, hardness-weighted contrastive learning further enhances its effectiveness, particularly with a 1.4-point improvement on out-of-distribution datasets, demonstrating the validity of our approach. Besides, we also compare the Focal-InfoNCE loss (ID 7), which also weights positive and negative pairs. Although it slightly improves performance on the out-of-distribution dataset, it reduces performance on the in-distribution dataset. As shown in Figure 1 and Table 1, our framework significantly increases the similarity gap between positive and negative pairs, thereby improving the model’s discriminative capability.
@@ 4.4 Zero-shot Video Retrieval
We also evaluate LLaVE on the widely used text-video retrieval datasets: MSR-VTT Xu et al. (2016) and MSVD Chen and Dolan (2011), to explore its generalization capability. There are two types of comparative models considered. The first type includes models trained on tens of millions of video-text data, such as InternVideo Wang et al. (2022), ViCLIP Wang et al. (2024a), UMT-L Li et al. (2023b), and InternVideo2-6B Wang et al. (2024b). The second type is completely zero-shot, trained only on text-image data using contrastive learning and directly evaluated on text-video data. Particularly, we focus on the two strongest models, VLM2Vec (LLaVA-OV-7B) and LamRA Liu et al. (2024), which is based on Qwen2-VL-7B and consists of two 7B models: LamRA-Ret and LamRA-Rank. LamRA-Ret retrieves the top-K candidates, while LamRA-Rank further re-ranks these retrieved candidates.
@@ 5 Related Work
As a significant research direction, multimodal embeddings aim to integrate the information from multiple modalities (e.g., vision and language) into a shared representation space, which enables seamless understanding across modalities. Early research primarily focuses on leveraging dual-stream architectures to separately encode texts and images. For instance, CLIP Radford et al. (2021), ALIGN Jia et al. (2021), BLIP Li et al. (2022), and SigLIP Zhai et al. (2023) all adopt dual-encoder frameworks, learning universal representations from large-scale weakly supervised image-text pairs through contrastive learning. To learn the more universal multimodal representations, UniIR Wei et al. (2024) proposes two fusion mechanisms to combine the visual and textual representations generated by the dual-encoder model. Although these models have achieved impressive results, they still face challenges in handling tasks such as interleaved image-text retrieval Wu et al. (2021) and instruction-following multimodal retrieval.
@@
To address the aforementioned issue, E5-V Jiang et al. (2024a) and VLM2Vec Jiang et al. (2024b) transform LMM into multimodal embedding models through contrastive learning, fully leveraging LMM’s powerful multimodal understanding capability and its inherent advantage in handling interleaved text-image input. Recently, a few concurrent studies further explored the application of LMMs in multimodal embeddings. For example, LamRA Liu et al. (2024) adopts the retrieval model to select the top-K candidates, which are then scored by the reranking models. Finally, the scores from the retrieval and reranking models are combined using a weighted sum to produce the final score for retrieval. MMRet Zhou et al. (2024) creates a large-scale multimodal instruction retrieval dataset called MegaPairs. By pretraining on this dataset, MMRet achieves SOTA results on MMEB.
@@
Negative samples play a crucial role in contrastive learning, Chen et al. (2020) show that incorporating more negative samples can enhance the performance of contrastive learning. Awasthi et al. (2022) further explore the impact of the number of negative samples from both theoretical and empirical perspectives. Moreover, Cai et al. (2020) demonstrate that hard negative samples are both necessary and sufficient to learn more discriminative representation, and Robinson et al. (2021) propose a hard negative sampling strategy where the user can control the hardness. The study most similar to ours is Focal-InfoNCE Hou and Li (2023), which weighs both positive and negative pairs based on their query-target similarities. Specifically, it uses a fixed threshold to determine the hardness of negative pairs, increasing the weight if the similarity exceeds the threshold and decreasing it otherwise. Unlike this work, our hardness-weighted contrastive learning introduces a reward model to dynamically estimate the hardness of negative pairs and applies weighting only to the negative pairs based on the estimated hardness. Notably, the reward model can be decoupled from the policy model.
@@ Appendix A Appendix
A.1 Hyperparameter Analysis
A.2 Training Details
@@
We present the training details of LLaVE in Table 5. Due to resource constraints, LLaVE-7B is trained on 910B GPUs, and the training time will significantly decrease when conducted on A100 GPUs. In addition, using more GPUs to adopt a larger batch size or higher resolution will improves the model’s performance.
A.3 Detailed results on MMEB

### v2

# arXiv:2503.04812v2[cs.CV] 01 Mar 2026

@@ 1 Introduction
Multimodal embedding models aim to encode inputs from any modality into vector representations, which then facilitate various multimodal tasks, such as image-text retrieval Wu et al. (2021a); Zhang et al. (2024a), automatic evaluation Hessel et al. (2021), and retrieval-augmented generation (RAG) Zhao et al. (2023). Although advanced pretrained vision-language models such as CLIP Radford et al. (2021), ALIGN Jia et al. (2021), and SigLIP Zhai et al. (2023) can provide unified representations for text and images, they face difficulties when dealing with more complex tasks. Particularly, they adopt a dual-encoder architecture that encodes images and text separately, leading to poor performance in tasks such as interleaved image-text retrieval Wu et al. (2021a).
In recent years, the rapid development and exceptional performance of large multimodal models (LMMs) have prompted researchers to focus increasingly on LMM-based multimodal embedding models. Compared to traditional pretrained vision-language models, LMMs not only demonstrate superior multimodal semantic understanding capabilities but also naturally support interleaved text-image inputs OpenAI (2023); Liu et al. (2023, 2024a); Lan et al. (2025); Hu et al. (2025). This advantage makes them more flexible and efficient in handling multimodal embedding tasks. As a representative work, Jiang et al. (2024b) construct the Massive Multimodal Embedding Benchmark (MMEB), which encompasses 4 meta-tasks and 36 datasets, and train multimodal embedding models based on LMMs. Experiment results demonstrate that by providing suitable task instructions and employing contrastive learning for training, LMM can significantly outperform existing multimodal embedding models and generalize effectively across diverse tasks.
@@ 2 Preliminary Study
SUN397 ImageNet-R DocVQA TextVQA CIRR Wiki-$S$-NQ MSCOCO RefCOCO-Matching
@@ 3 Our Framework
In this section, we first illustrate the inherent consistency between preference learning and contrastive learning. We then propose a simple yet effective framework that incorporates hardness-weighted contrastive learning (See Section 3.1) and cross-device negative sample gathering (See Section 3).
@@ 3.1 Hardness-Weighted Contrastive Learning
$\mathcal{L}_{i}=-\log\frac{e^{{r_{\pi}(q_{i},t_{i})}}}{e^{{r_{\pi}(q_{i},t_{i})}}+\sum_{j\neq i}^{N}w_{ij}\cdot e^{{{r_{\pi}(q_{i},t_{j})}}}},$ (4)
where $w_{ij}$ represents the weight of learning difficulty. As shown in Figure 2, to estimate the learning difficulty of pairs, we introduce a reward model $r_{\theta}$ and set $w_{ij}=e^{r_{\theta}(q_{i},t_{j})}$. Accordingly, the policy model adjusts its learning of different negative pairs based on the feedback from the reward model. In this work, to achieve higher training efficiency and simpler implementation, we update the reward model $r_{\theta}$ to keep aligned with the policy model $r_{\pi}$ after each step. The reward model does not perform backpropagation, i.e., $r_{\theta}(q_{i},t_{j})$ = $\alpha\cdot\mathrm{sg}(s_{ij})$, where $\alpha$ is a hyperparameter and $\mathrm{sg}(\cdot)$ denotes the stop-gradient operation. When the reward model is set to be the same as the policy model, another advantage is that assigning a higher reward to a negative sample indicates that the policy model finds it more difficult to distinguish, thereby enhancing its learning on currently hard samples. Note that $r_{\theta}$ can also adopt model structures other than the policy model. Finally, $\mathcal{L}_{i}$ is defined as
@@
From Equation 6, we can observe that the gradients of the negative pairs are proportional to the product $r_{\theta}(q_{i},t_{j})$. which implies that the greater the learning difficulty of a negative pair, the more significant its role in the gradient update.
@@ 3.2 Cross-Device Negative Sample Gathering
Figure 3: An example of cross-device negative sample gathering (NN=44 and KK=33). The plus signs represent positive pairs, and the minus signs represent negative pairs. Each device calculates the similarity between its own queries and the targets on all other devices, which is then used for loss computation.
@@ 4.1 Setup
Table 2: Results on the MMEB benchmark. IND represents the in-distribution dataset, and OOD represents the out-of-distribution dataset. In UniIR, the $F$ and SF subscripts under CLIP or BLIP represent feature-level fusion and score-level fusion, respectively. LLaVA-NeXT-7B-LR indicates the use of low-resolution (336×336) image inputs, while LLaVA-NeXT-7B-HR refers to the use of high-resolution (1344×1344) image inputs. The reported scores are the average Precision@1 over the corresponding datasets. The best results are marked in bold, and the second-best results are underlined. Part of the baseline results are sourced from Jiang et al. (2024b) and Zhou et al. (2024). Detailed results and qualitative evaluations can be found in Appendix A.6 and Section 4.5.
@@
Our trained model, LLaVE, includes three scales: 0.5B, 2B, and 7B, based on LLaVA-OV-0.5B Li et al. (2024), Aquila-VL-2B Gu et al. (2024), and LLaVA-OV-7B Li et al. (2024), respectively. To facilitate community use, the training code for LLaVE is built on the widely-used Transformers Wolf et al. (2020) and DeepSpeed Rasley et al. (2020) packages. We use a batch size of 256 by gradient accumulation, set the weighting hyperparameter $\alpha$ to 9  [^1] We empirically analyze the impact of $\alpha$ on model performance in Appendix A.4., and impose a total length limit of 4096. Furthermore, we employ the Higher Anyres technique Li et al. (2024) to support high-resolution images, setting the maximum image resolution to 672 $\times$ 672. The learning rate is set to 1e-5 for LLaVE-0.5B and LLaVE-2B, and 5e-6 for LLaVE-7B. For efficient training, we freeze the vision encoder and train the model for one epoch using the DeepSpeed ZeRO-3 strategy. Regarding training costs, LLaVE-0.5B and LLaVE-2B are trained on 8 NVIDIA A100 GPUs (40GB) for 12 and 17 hours, respectively, while LLaVE-7B is trained on 16 Ascend 910B GPUs (64GB) for 33 hours. More details can be found in Appendix A.1.
@@ 4.2 Main Results
Although previous models have achieved strong performance, our LLaVE series demonstrates consistent improvements over the best baseline across all metrics. LLaVE-7B achieves a state-of-the-art overall score of 70.3, outperforming the previous SOTA model, MMRet, by 6.2 points and surpassing VLM2Vec (LLaVA-OV-7B) by 4.5 points. In grounding, LLaVE-7B attains an impressive score of 91.9, a +4.6 point improvement over VLM2Vec. LLaVE-7B also leads in VQA (65.4) and classification (65.7), with improvements of +4.3 and +2.2 points, respectively. In addition, we observe that the performance of LLaVE models scales consistently with model size, indicating that our framework has excellent scalability. It is worth mentioning that the performance of LLaVE-0.5B is already comparable to VLM2Vec (Phi-3.5-V-4B), while LLaVE-2B achieves an overall score of 65.2, surpassing the previously pretrained MMRet-7B that utilizes an additional 27M image-text pairs.
@@ 4.3 Ablation Study
Although the model (ID 6) already achieves near-SOTA performance, hardness-weighted contrastive learning further enhances its effectiveness, particularly with a 1.4-point improvement on out-of-distribution datasets, demonstrating its complementarity. Besides, we also compare the Focal-InfoNCE loss (ID 7), which also weights positive and negative pairs. Although it slightly improves performance on the out-of-distribution dataset, it reduces performance on the in-distribution dataset. As shown in Figure 1 and Table 1, our framework significantly increases the similarity gap between positive and negative pairs, thereby improving the model’s discriminative capability.
@@ 4.4 Zero-shot Video Retrieval
We also evaluate LLaVE on the widely used text-video retrieval datasets: MSR-VTT Xu et al. (2016) and MSVD Chen and Dolan (2011), to explore its generalization capability. There are two types of comparative models considered. The first type includes models trained on tens of millions of video-text data, such as InternVideo Wang et al. (2022), ViCLIP Wang et al. (2024a), UMT-L Li et al. (2023b), and InternVideo2-6B Wang et al. (2024b). The second type is completely zero-shot, trained only on text-image data using contrastive learning and directly evaluated on text-video data. Particularly, we focus on the two strongest models, VLM2Vec (LLaVA-OV-7B) and LamRA Liu et al. (2024b), which is based on Qwen2-VL-7B and consists of two 7B models: LamRA-Ret and LamRA-Rank. LamRA-Ret retrieves the top-K candidates, while LamRA-Rank further re-ranks these retrieved candidates.
@@ 5 Related Work
As a significant research direction, multimodal embeddings aim to integrate the information from multiple modalities (e.g., vision and language) into a shared representation space, which enables seamless understanding across modalities. Early research primarily focuses on leveraging dual-stream architectures to separately encode texts and images. For instance, CLIP Radford et al. (2021), ALIGN Jia et al. (2021), BLIP Li et al. (2022), and SigLIP Zhai et al. (2023) all adopt dual-encoder frameworks, learning universal representations from large-scale weakly supervised image-text pairs through contrastive learning. To learn the more universal multimodal representations, UniIR Wei et al. (2024) proposes two fusion mechanisms to combine the visual and textual representations generated by the dual-encoder model. Although these models have achieved impressive results, they still face challenges in handling tasks such as interleaved image-text retrieval Wu et al. (2021a) and instruction-following multimodal retrieval.
@@
To address the aforementioned issue, E5-V Jiang et al. (2024a) and VLM2Vec Jiang et al. (2024b) transform LMM into multimodal embedding models through contrastive learning, fully leveraging LMM’s powerful multimodal understanding capability and its inherent advantage in handling interleaved text-image input. Recently, a few concurrent studies Zhang et al. (2025); Gu et al. (2025); Thirukovalluru et al. (2025) further explored the application of LMMs in multimodal embeddings. For example, LamRA Liu et al. (2024b) adopts the retrieval model to select the top-K candidates, which are then scored by the reranking models. Finally, the scores from the retrieval and reranking models are combined using a weighted sum to produce the final score for retrieval. MMRet Zhou et al. (2024) creates a large-scale multimodal instruction retrieval dataset called MegaPairs. By pretraining on this dataset, MMRet achieves SOTA results on MMEB.
@@
Contrastive learning enables models to learn effective representations by distinguishing between positive and negative samples, and it has been widely applied across various domains Kipf et al. (2020); Wu et al. (2021b); Lan et al. (2023); Yin et al. (2023); Zhang et al. (2024c). Negative samples play a crucial role in contrastive learning, Chen et al. (2020) show that incorporating more negative samples can enhance the performance of contrastive learning. Awasthi et al. (2022) further explore the impact of the number of negative samples from both theoretical and empirical perspectives. Moreover, Cai et al. (2020) demonstrate that hard negative samples are both necessary and sufficient to learn more discriminative representation, and Robinson et al. (2021) propose a hard negative sampling strategy where the user can control the hardness. The study most similar to ours is Focal-InfoNCE Hou and Li (2023), which weighs both positive and negative pairs based on their query-target similarities. Specifically, it uses a fixed threshold to determine the hardness of negative pairs, increasing the weight if the similarity exceeds the threshold and decreasing it otherwise. Unlike this work, our hardness-weighted contrastive learning introduces a reward model to dynamically estimate the hardness of negative pairs and applies weighting only to the negative pairs based on the estimated hardness. Notably, the reward model can be decoupled from the policy model.
@@ 6 Conclusion
Acknowledgments
The project was supported by National Key R&D Program of China (No. 2022ZD0160501), Natural Science Foundation of Fujian Province of China (No. 2024J011001), and the Public Technology Service Platform Project of Xiamen (No.3502Z20231043). We also thank the reviewers for their insightful comments.
@@ Appendix A Appendix
A.1 Training Details
@@
We present the training details of LLaVE in Table 5. To save memory and accelerate training, we adopt Gradient Checkpointing and Flash Attention Dao et al. (2022) techniques. Due to resource constraints, LLaVE-7B is trained on 910B GPUs, and the training time will significantly decrease when conducted on A100 GPUs. In addition, using more GPUs to adopt a larger batch size or higher resolution will improve the model’s performance.
A.2 The Impact of Different Batch Sizes
Figure 5: The impact of batch size on overall model performance when training for one epoch.
As shown in Figure 5, we also explore the impact of different batch sizes on the overall performance of the model when training for one epoch. It can be observed that as the batch size increases, the overall performance of the model improves. However, since the total number of trained epochs does not increase, larger batch sizes do not achieve further performance improvements. This is similar to the trend observed by Zhai et al. (2023), as large batch sizes need a sufficiently long schedule to ramp up. Besides, we find that the similarity distribution does not differ significantly across different batch sizes, which means that the similarity distribution remains relatively stable using different batch sizes.
Figure 6: Impact of training epochs on IND and OOD performance.
A.3 The Impact of Training Epochs
To further investigate the impact of training duration, we fine-tune LLaVE-2B for up to five epochs. The results in Figure 6 show that increasing the number of epochs initially enhances in-domain performance and overall accuracy. However, after three epochs, overfitting appears, which degrades out-of-domain performance and limits further overall improvements. For the sake of training efficiency and generalization, we therefore train for one epoch throughout this paper.
Model Per Meta-Task Score Average Score
Classification VQA Retrieval Grounding IND OOD Overall
# Datasets 10 10 12 4 20 16 36
VLM2Vec (Qwen2-VL-2B) 59.0 49.4 65.4 73.4 66.0 52.6 60.1
VLM2Vec (Qwen2-VL-7B) 62.6 57.8 69.9 81.7 72.2 57.8 65.8
LLaVE-2B (Qwen2-VL-2B) 64.3 58.5 66.5 76.9 71.1 57.1 64.8
LLaVE-7B (Qwen2-VL-7B) 65.4 65.2 70.6 84.8 74.8 62.3 69.2
Table 6: Comparison of LLaVE and VLM2Vec based on different LMMs.
A.4 Hyperparameter Analysis
A.5 Generalization of LLaVE Across LMMs
To further validate the generality of our proposed methods, we extend LLaVE to other widely used LMMs and evaluated it on MMEB. Specifically, we apply our methods to Qwen2-VL models of different scales (2B and 7B) and employ the same training settings as VLM2Vec, including identical batch sizes and training steps. To further validate the generalizability of our proposed methods, we extend LLaVE to other widely used LMMs and evaluate its performance on MMEB. Specifically, we apply our methods to Qwen2-VL models of varying scales (2B and 7B) and employ the same training settings as VLM2Vec. As shown in Table 5, when using the same LMM as the backbone, LLaVE still significantly outperforms VLM2Vec. In particular, LLaVE-2B achieves an improvement of 4.7 over VLM2Vec-2B, and LLaVE-7B achieves an improvement of 3.4 over VLM2Vec-7B. These results fully demonstrate the strong generalization capability of our method.
A.6 Detailed results on MMEB